OpenStudy (anonymous):

Solve the following system of equations: 3x - 2y = 6 6x - 4y = 14 (0, 0) (6, 14) Infinitely many solutions No solutions

OpenStudy (anonymous):

@YesThisIsDog , @brittanydosey , @phi , @snazzychazzy , @farmgirl11 , @horsegirl325

OpenStudy (anonymous):

Ask @iGreen

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

If you input the answers you will find that it is D but igreen might answer different

OpenStudy (igreen):

Let's rearrange to slope-intercept form. 3x - 2y = 6 Subtract 3x to both sides: -2y = -3x + 6 Divide -2 to both sides: y = 3/2x + 6/-2 What's 6 / -2?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

I have no idea

OpenStudy (igreen):

Yes, so we have: y = 3/2x - 3 6x - 4y = 14 Subtract 6x to both sides: -4y = -6x + 14 Divide -4 to both sides: y = 6/4x + 14/-4 What's 14 / -4?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

I'm not sure how 14 and 4 go into each other.

OpenStudy (igreen):

14 / -4 = -3.5 So we have: y = 6/4x - 3.5 Simplify 6/4. 6 / 2 = ? 4 / 2 = ?

OpenStudy (anonymous):

6/2 = 3 4/2 = 2

OpenStudy (igreen):

Yes, so we have: y = 3/2x - 3.5 and y = 3/2x - 3 Therefore the line's are parallel and have no solution.

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

can i have medal for answering first?